BankruptTiaa:

Olivia needs $36 to buy a coat. She has saved $8 and plans to work as a babysitter to earn $7 per hour. Which inequality shows the minimum number of hours, n, that Olivia should work as a babysitter to earn enough to buy the coat? 8 + 7n ≤ 36, so n ≤ 4 8 + 7n ≥ 36, so n ≥ 4 7n ≥ 36 + 8, so n ≥ 6.3 7n ≤ 36 + 8, so n ≤ 6.3

BankruptTiaa:

help meee

Florisalreadytaken:

Well, Olivia already has got her $8, so we know that the formula will have \( 8+ \ ... \) if she doesnt have AT LEAST \(\geq\) $36, then she cannot get the coat. Therefore the right answer is \( 8 + 7n \geq 36 \) where she must work AT LEAST 4 hours \( n \geq 4 \)

BankruptTiaa:

@florisalreadytaken wrote:
Well, Olivia already has got her $8, so we know that the formula will have \( 8+ \ ... \) if she doesnt have AT LEAST \(\geq\) $36, then she cannot get the coat. Therefore the right answer is \( 8 + 7n \geq 36 \) where she must work AT LEAST 4 hours \( n \geq 4 \)
Ty